# VeriLLM: A Lightweight Framework for Publicly Verifiable Decentralized Inference

### v3

@@ 2.2 Design Goals
G1. Public Verifiability. Any participant can verify inference correctness without relying on trust or replication. All hidden states and token outputs are bound by cryptographic commitments and, through a commit–then–sample procedure [sun2024zkllm], any participant can open selected entries and check them against the model specification. This goal directly mitigates C1 and C2 by allowing verifiable detection of computation deviation or lazy verification. The verification process is auditable both off-chain (via hidden-state recomputation) and on-chain (via sampled scalar checks).
G2. Low Overhead. Verification overhead remains below 1% of the total inference cost. The scheme achieves this by limiting on-chain operations to sampled scalar openings rather than full tensor uploads, and by parallelizing verifier recomputation across layer segments [rajbhandari2022deep]. This goal ensures practical deployability under realistic hardware and network conditions, addressing the performance constraints implied by C1 (economic motivation to cheat).
G3. One-Honest-Verifier Security. Correctness is guaranteed as long as at least one verifier executes the protocol honestly. Even if all but one verifier collude with the Inferencer or Scheduler, the presence of a single honest verifier ensures that inconsistencies or forged hidden states will be revealed either through the on-chain sampling phase or via zero-knowledge dispute proofs [qu2025zkgpt, chen2024zkml]. This property directly counters C2 and C3 by removing the need for a majority or trusted quorum.
G4. Deterministic Accountability. Every message and hidden-state relay is traceable to a verifiable, signed commitment. Each segment boundary state is committed via a Merkle root [merkle1989certified] and signed by the responsible node. These signatures are registered on-chain, ensuring that any alteration, omission, or forgery by a malicious Scheduler (C3) or Inferencer (C1) is cryptographically detectable. This binding of state to identity guarantees tamper-evident trace integrity.
G5. Task-Type Indistinguishability. Workers cannot distinguish whether a job is an inference or a verification task. Inferencers and Verifiers receive identical message formats and API calls from the Scheduler, preventing selective deviation based on task type (mitigating C4). A node that attempts to skip verification will be caught with high probability during unpredictable on-chain sampling [yuan2022decentralized].
G6. Compatibility with Decentralized Deployment. The protocol operates robustly across heterogeneous GPUs and compute environments. VeriLLM tolerates bounded floating-point divergence by using calibrated statistical thresholds in hidden-state comparison [shoeybi2019megatron]. This ensures that cross-device verification remains feasible and fair while preventing adversaries from exploiting hardware-level drift to mask malicious changes.
2.3 Threat–Goal Alignment
Table I summarizes the alignment between adversarial capabilities and the corresponding design goals that mitigate each threat.
TABLE I: Threat–Goal Alignment Matrix for VeriLLM.
@@ 2.3 Threat–Goal Alignment
The combination of these goals forms a security envelope suited for large-scale decentralized inference. Public verifiability (G1) and deterministic accountability (G4) guarantee that every computation step is externally auditable. Low-overhead design (G2) ensures practical adoption without compromising latency or throughput. The one-honest-verifier guarantee (G3) provides strong correctness in adversarial environments, while task-type indistinguishability (G5) closes adaptive attack channels. Finally, cross-device compatibility (G6) grounds the protocol in realistic heterogeneous hardware deployments [rajbhandari2022deep, shoeybi2019megatron]. Together, these principles ensure that VeriLLM achieves transparent, efficient, and verifiable LLM inference under adversarial conditions typical of open, decentralized systems.
3 Related Work
3.1 Activation Sampling & Compact Commitment
Lightweight methods verify inference integrity by selectively committing to critical activation states. TOPLOC/INTELLECT [ong2025toploc] commits to the top-128 final-layer activations every 32 tokens, compressing data 1000× via polynomial congruence (258 bytes per commitment). However, its verification process requires a trusted third party to re-run the computation and provide a ground truth, which reintroduces a central trust assumption. TensorBlock [proof-of-cache-github] extends this to KV-cache sampling using deterministic token/layer selection, enabling validators to recompute partial trajectories from checkpoints. VeriSplit [zhang2024verisplit] supports private outsourcing through linear operator masking: clients add noise to inputs, workers return encrypted activations, and clients locally denoise results using precomputed terms, with Merkle trees enabling partial verification. However, their probabilistic security cannot guarantee the detection of sophisticated model tampering, and input privacy relies on noise injection techniques that are vulnerable to advanced reconstruction attacks.
3.2 Zero-Knowledge Machine Learning (ZKML)
Sumcheck-based Protocols: zkCNN [liu2021zkcnn] accelerates convolution verification by transforming operations to the frequency domain. zkLLM [sun2024zkllm] reduces communication by 90% for non-arithmetic operations (e.g., Softmax) via the tlookup protocol and exploits translation invariance. QAP-based Protocols: vCNN [lee2024vcnn] represents convolutions as Quadratic Polynomial Programs (QPP), reducing multiplicative gates exponentially. pvCNN [weng2023pvcnn] further optimizes tensor operations via Quadratic Matrix Programs (QMP) and batch verification. ZKML [chen2024zkml] offers modular circuit design with 43 pre-built layer circuits and automated optimization. zkGPT [qu2025zkgpt] leverages multi-thread parallelization for LLM inference-proof acceleration. Lookup arguments [campanelli2024lookup] enable matrix-level privacy with 30% faster proofs. VOLE-based Protocols: Mystique [weng2021mystique] achieves sublinear communication for matrix multiplication using sVOLE-based tensor compression. Recent frameworks [hao2024scalable] optimize nonlinear functions via table lookups and digital decomposition, reducing constraints by over 99%.
System & Hardware Optimization: zkLoRA [roy2025zklora] supports verifiable fine-tuning via incremental proofs for LoRA modules. EZKL [ezkl-github] automates ZKP circuit generation from ONNX models. DeepProve [deep-prove-github] accelerates GPU-based ZKML (54–158× faster proofs via parallelized SNARKs). Sertn [sertn-avs-github] employs optimistic ”Staked Deferred Proofs” with economic slashing for invalid results. However, these methods impose prohibitive prover costs (hours per inference), require trusted setups, and suffer from rigidity in supporting novel architectures due to circuit specialization.
3.3 Trusted Execution Environments (TEEs)
Atoma Network [atoma-infer-github] combines TEE attestation with sampling-based consensus for deterministic execution. evML [evml-github] uses TEE-based remote attestation (e.g., Google/Apple frameworks) with probabilistic auditing. nesa.ai [nesa-github] orchestrates heterogeneous TEEs for secure multi-party inference. Phala Network [phala2024] extends this paradigm by providing a ”Confidential AI” cloud platform that enables standard Dockerized AI applications to run within GPU-equipped TEEs. However, they inherit vulnerabilities to side-channel exploits (e.g., cache-timing attacks), centralize trust in proprietary hardware (Intel SGX/ARM TrustZone), and exhibit limited auditability across heterogeneous environments.
3.4 Consensus & Optimistic Approaches
Verde [arun2025verde] employs optimistic dispute resolution, recomputing only divergent operators (e.g., single attention heads) to minimize overhead. Mira [mira-network] leverages multi-model consensus against hallucinations using domain-specific thresholds. Ambient [ambient-ai] introduces Proof-of-Logits (PoL), hashing intermediate logits for spot-check verification at $0.1\%$ overhead. However, they introduce delayed finality during dispute windows and implicitly assume honest majority participation.
4 Background & Problem Statements
@@
Full-Sequence Prefill for Parallel Verification. Autoregressive large language models (LLMs) rely on causal (left-to-right) self-attention, in which each hidden state $h_{t}$ depends on all preceding tokens $\{x_{1},\dots,x_{t-1}\}$. This dependency constrains the Decoder to operate sequentially, forming the dominant computational bottleneck in inference. Practical serving systems mitigate this through a Prefill phase that computes the contextual representations for all prompt tokens in parallel and materializes the key–value (KV) cache, allowing subsequent token-by-token decoding to reuse these cached features [agrawal2023, shazeer2019fast, rajbhandari2022deep]. At verification time, this architectural separation becomes even more advantageous: since the verifier observes the full input–output sequence $(x_{1},\dots,x_{T})$, it can execute a single forward pass under the causal mask to recover all hidden states simultaneously, without performing the autoregressive loop.
@@
$H^{\prime}=\mathcal{F}_{\theta}(X_{\text{prompt}}\|X_{\text{output}}),$ (1)
@@
$\frac{C_{\text{verify}}}{C_{\text{infer}}}=\frac{T_{\text{prompt}}}{T_{\text{prompt}}+T_{\text{output}}}=0.01,$ (2)
@@
Lazy Verification. A naive verification baseline re-executes the entire model forward pass under identical hyperparameters and compares the resulting hidden states with those reported by the Inferencer. However, because honest inference is the common case, verifiers have an incentive to skip computation and submit a trivial True verdict—an instance of the lazy-verification problem, aka. the verifier’s dilemma [luu2015demystifying]. Such behavior gradually undermines system integrity and leads to unverifiable results.
@@
To address this, VeriLLM employs a lightweight commit–sample–check protocol. Each verifier $v_{i}$ commits to its hidden states by posting a Merkle root $r_{i}=\mathsf{Merkle}(H_{i})$ [merkle1989certified]. After commitments are finalized, a smart contract derives random challenge indices $\mathcal{S}=\{s_{1},\dots,s_{k}\}$ from a verifiable random function (VRF) [micali1999verifiable], ensuring unpredictability and fairness. Each verifier must open the corresponding leaves $\{H_{i}[s_{j}]\}_{j=1}^{k}$ and submit inclusion proofs $\pi_{i}[s_{j}]$ to the contract. Let $\delta(H_{i}[s_{j}],H^{*}[s_{j}])$ denote the numerical deviation from the Inferencer’s committed states. The contract enforces:
@@
$\mathsf{Slash}(v_{i})=\begin{cases}1,&\text{if }\exists j\text{ s.t. }\delta(H_{i}[s_{j}],H^{*}[s_{j}])>\epsilon,\\ 0,&\text{otherwise,}\end{cases}$ (3)
@@
where $\epsilon$ is a tolerance threshold capturing floating-point nondeterminism [dettmers2022llmint8, frantar2023gptq]. Verifiers whose openings align with the Inferencer’s states within $\epsilon$ receive rewards, while inconsistent openings trigger slashing. To ensure continuous participation, a peer-consistency mechanism statistically compares verifier outputs and rewards correlation with the empirical distribution expected from honest computation, thus discouraging collusion and inactive nodes.
@@
Security under Sampling Majority. Conventional consensus protocols [castro1999practical, cachin2011introduction] rely on an honest-majority assumption to guarantee safety and liveness. In contrast, VeriLLM achieves correctness under the weaker one-honest-verifier (OHV) assumption. Let $n$ denote the total number of verifiers, $f$ the number of malicious ones, and $k$ the committee size for sampling-based verification. The probability that all $k$ selected verifiers are malicious is:
$P_{\text{fail}}=\frac{\binom{f}{k}}{\binom{n}{k}}$. (4)
For a fixed corruption rate $\rho=f/n<0.5$, $P_{\text{fail}}$ decreases exponentially in $k$:
$P_{\text{fail}}\sim\rho^{k},$ (5)
@@
which becomes negligible even for modest committee sizes (e.g., $k=10$). Thus, VeriLLM maintains strong probabilistic safety guarantees without global replication. To further reduce risk under targeted committee attacks, the system employs a dynamic verifier-selection mechanism that re-samples verifiers over epochs, ensuring that adversarial control of a fixed subset cannot persist across rounds. Together with a re-verification fallback procedure, this mechanism ensures that at least one honest verifier detects and reports inconsistencies with overwhelming probability.
5 Protocol Design
@@
Figure 1: System architecture of VeriLLM. The scheduler employs a verifiable random function (VRF) for unbiased node selection and hidden-state sampling. Each node group performs prefill and decoding, commits hidden states, and submits proofs to the blockchain, which verifies results, counts votes, and applies reward or penalty mechanisms to ensure verifiable decentralized inference.
5.1 System Overview
@@
VeriLLM is a publicly verifiable, decentralized inference framework for large language models (LLMs) built atop a blockchain substrate. The system aims to achieve three goals simultaneously: (i) correctness that is publicly auditable, (ii) low verification overhead compatible with real-time inference, and (iii) transparent incentive enforcement through smart contracts. To this end, VeriLLM comprises three core components: a scheduler, a set of homogeneous node groups, and a family of on-chain verification contracts. These components collectively coordinate the inference, verification, and reward mechanisms to ensure that every published result is both correct and economically accountable.
@@
  • Scheduler. The scheduler orchestrates task placement, resource allocation, and verification planning. For each inference job, it selects a candidate set of node groups capable of jointly executing the model’s layers or tensor partitions. Within each selected group, the scheduler designates one Inferencer and several Verifiers to handle computation and validation, respectively. Although logically centralized, the scheduler’s randomness and decisions are publicly verifiable: both role assignment and checkpoint sampling derive from a verifiable random function (VRF) [micali1999verifiable], producing unpredictable yet reproducible outcomes. This prevents adversaries from precomputing which nodes will be responsible for verification or from biasing verifier selection. The scheduler also records the mapping of job identifiers, selected groups, and Merkle root commitments of intermediate states onto the blockchain for immutable traceability.
  • Node Groups. VeriLLM organizes computation across groups of homogeneous GPU workers to simplify performance balancing and state consistency. Each node group corresponds to one or more identical Transformer layers, including all weights, normalization statistics, and kernel parameters. This design aligns with modern distributed inference frameworks such as tensor- and pipeline-parallel architectures [rajbhandari2022deep, shoeybi2019megatron]. Within a group, a node can serve either as an Inferencer or a Verifier depending on the scheduler’s role assignment for the current job. To mitigate role-conditioned adversarial strategies, both roles share a unified execution interface: from the node’s perspective, the input tensor, KV-cache access pattern, and runtime sequence length appear identical for inference and verification. Consequently, at the start of the Prefill phase, a worker cannot distinguish whether it is performing live inference or verifying another node’s computation. This task-type indistinguishability provides a built-in security property that limits strategic deviation, since any deviation risks immediate detection if the task later turns out to be a verification probe.
@@
  • Verification Contracts. The verification layer is implemented through on-chain smart contracts [buterin2014next, wood2014ethereum]. These contracts collect verifier submissions, validate sampling correctness, and finalize verification outcomes through contract-level consensus. Each participating node is required to deposit a stake when registering for inference or verification. The contracts automatically enforce rewards and penalties based on the outcome of verification, slashing misbehaving nodes when inconsistencies or invalid openings are detected. To minimize gas costs, VeriLLM adopts a two-tier protocol: (i) the heavy-lift computation and Merkle proof generation occur off-chain; (ii) only small digests, VRF proofs, and sampled Merkle openings are submitted on-chain. This hybrid design maintains full public verifiability while keeping on-chain cost sublinear in model size. Moreover, the contract interface exposes an event log of all verification transactions, allowing any third-party observer to independently re-verify computations, thereby realizing trustless auditability.
Overall, the VeriLLM system model integrates randomness, cryptographic accountability, and economic incentives within a unified architecture. The scheduler enforces randomness and non-manipulability, the node groups ensure hardware-level symmetry and indistinguishability, and the contracts provide immutable verification and incentive enforcement. Together, these elements enable decentralized inference to achieve correctness and efficiency comparable to centralized inference systems, while remaining fully open and publicly verifiable. End-to-end workflow. (1) A client submits an inference request to the scheduler, which first selects the node groups that can jointly serve the job and then, using a VRF, chooses one Inferencer and several Verifiers from each group. (2) The Inferencer executes the forward pass in a pipelined manner, streaming intermediate hidden states to the scheduler; the scheduler records these hidden states and forwards them among inference nodes as needed. (3) After inference completes, the scheduler dispatches the corresponding initial hidden states to the assigned Verifiers. Each Verifier runs a prefill computation to reproduce the target hidden states, compares them with the Inferencer’s hidden states, and submits the comparison result to the smart contract on-chain. (4) To deter “lazy” verification, the scheduler samples hidden-state indices; Verifiers must submit the sampled hidden states for on-chain comparison, and the contract rewards or penalizes Verifiers according to the on-chain outcome.
5.2 Registration

### v4

@@ 3.2. Design Goals
G1. Public Verifiability. Any participant can verify inference correctness without relying on trust or replication. All hidden states and token outputs are bound by cryptographic commitments and, through a commit–then–sample procedure (Sun et al., 2024), any participant can open selected entries and check them against the model specification. This goal directly mitigates C1 and C2 by allowing verifiable detection of computation deviation or lazy verification. The verification process is auditable both off-chain (via hidden-state recomputation) and on-chain (via sampled scalar checks).
G2. Low Overhead. Verification overhead remains below 1% of the total inference cost. The scheme achieves this by limiting on-chain operations to sampled scalar openings rather than full tensor uploads, and by parallelizing verifier recomputation across layer segments (Rajbhandari et al., 2022). This goal ensures practical deployability under realistic hardware and network conditions, addressing the performance constraints implied by C1 (economic motivation to cheat).
G3. Deterministic Accountability. Every message and hidden-state relay is traceable to a verifiable, signed commitment. Each segment boundary state is committed via a Merkle root (Merkle, 1989) and signed by the responsible node. These signatures are registered on-chain, ensuring that any alteration, omission, or forgery by a malicious Scheduler (C3) or Inferencer (C1) is cryptographically detectable. This binding of state to identity guarantees tamper-evident trace integrity.
G4. Task-Type Indistinguishability. Workers cannot distinguish whether a job is an inference or a verification task. Inferencers and Verifiers receive identical message formats and API calls from the Scheduler, preventing selective deviation based on task type (mitigating C4). A node that attempts to skip verification will be caught with high probability during unpredictable on-chain sampling (Yuan et al., 2022).
G5. Compatibility with Decentralized Deployment. The protocol operates robustly across heterogeneous GPUs and compute environments. VeriLLM tolerates bounded floating-point divergence by using calibrated statistical thresholds in hidden-state comparison (Shoeybi et al., 2019). This ensures that cross-device verification remains feasible and fair while preventing adversaries from exploiting hardware-level drift to mask malicious changes.
3.3. Threat–Goal Alignment
Table 1 summarizes the alignment between adversarial capabilities and the corresponding design goals that mitigate each threat.
Table 1. Threat–Goal Alignment Matrix for VeriLLM.
@@ 3.3. Threat–Goal Alignment
The combination of these goals forms a security envelope suited for large-scale decentralized inference. Public verifiability (G1) and deterministic accountability (G3) guarantee that every computation step is externally auditable. Low-overhead design (G2) ensures practical adoption without compromising latency or throughput. The Deterministic Accountability guarantee (G3) provides strong correctness in adversarial environments, while task-type indistinguishability (G4) closes adaptive attack channels. Finally, cross-device compatibility (G5) grounds the protocol in realistic heterogeneous hardware deployments (Rajbhandari et al., 2022; Shoeybi et al., 2019). Together, these principles ensure that VeriLLM achieves transparent, efficient, and verifiable LLM inference under adversarial conditions typical of open, decentralized systems.
4. Background & Problem Statements
@@
Full-Sequence Prefill for Parallel Verification. Autoregressive large language models (LLMs) rely on causal (left-to-right) self-attention, in which each hidden state $h_{t}$ depends on all preceding tokens $\{x_{1},\dots,x_{t-1}\}$. This dependency constrains the Decoder to operate sequentially, forming the dominant computational bottleneck in inference. Practical serving systems mitigate this through a Prefill phase that computes the contextual representations for all prompt tokens in parallel and materializes the key–value (KV) cache, allowing subsequent token-by-token decoding to reuse these cached features (Agrawal et al., 2023; Shazeer, 2019; Rajbhandari et al., 2022). At verification time, this architectural separation becomes even more advantageous: since the verifier observes the full input–output sequence $(x_{1},\dots,x_{T})$, it can execute a single forward pass under the causal mask to recover all hidden states simultaneously, without performing the autoregressive loop.
@@
(1) $H^{\prime}=\mathcal{F}_{\theta}(X_{\text{prompt}}\|X_{\text{output}}),$
@@
(2) $\frac{C_{\text{verify}}}{C_{\text{infer}}}=\frac{T_{\text{prompt}}}{T_{\text{prompt}}+T_{\text{output}}}=0.01,$
@@
Lazy Verification. A naive verification baseline re-executes the entire model forward pass under identical hyperparameters and compares the resulting hidden states with those reported by the Inferencer. However, because honest inference is the common case, verifiers have an incentive to skip computation and submit a trivial True verdict—an instance of the lazy-verification problem, aka. the verifier’s dilemma (Luu et al., 2015; Zhao et al., 2024). Such behavior gradually undermines system integrity and leads to unverifiable results.
@@
To address this, VeriLLM employs a lightweight commit–sample–check protocol. Each verifier $v_{i}$ commits to its hidden states by posting a Merkle root $r_{i}=\mathsf{Merkle}(H_{i})$ (Merkle, 1989). After commitments are finalized, a smart contract derives random challenge indices $\mathcal{S}=\{s_{1},\dots,s_{k}\}$ from a verifiable random function (VRF) (Micali et al., 1999), ensuring unpredictability and fairness. Each verifier must open the corresponding leaves $\{H_{i}[s_{j}]\}_{j=1}^{k}$ and submit inclusion proofs $\pi_{i}[s_{j}]$ to the contract. Let $\delta(H_{i}[s_{j}],H^{*}[s_{j}])$ denote the numerical deviation from the Inferencer’s committed states. The contract enforces:
@@
(3) $\mathsf{Slash}(v_{i})=\begin{cases}1,&\text{if }\exists j\text{ s.t. }\delta(H_{i}[s_{j}],H^{*}[s_{j}])>\epsilon,\\ 0,&\text{otherwise,}\end{cases}$
@@
where $\epsilon$ is a tolerance threshold capturing floating-point nondeterminism (Dettmers et al., 2022; Frantar and Alistarh, 2023). Verifiers whose openings align with the Inferencer’s states within $\epsilon$ receive rewards, while inconsistent openings trigger slashing. To ensure continuous participation, a peer-consistency mechanism statistically compares verifier outputs and rewards correlation with the empirical distribution expected from honest computation, thus discouraging collusion and inactive nodes.
@@
Security under Global Honest Majority. Conventional consensus protocols (Castro and Liskov, 1999; Cachin et al., 2011) rely on an honest-majority assumption among sampled nodes to guarantee safety and liveness. In contrast, VeriLLM achieves correctness under a more realistic global honest majority assumption. Let $n$ denote the total number of verifiers, $f$ the number of malicious ones, $k$ the committee size for sampling-based verification, and we assume $k$ is odd. Then in conventional protocols, the security fails if and only if more than $\frac{k}{2}$ committee nodes are malicious. Hence, we have:
(4) $\tilde{P}_{\text{fail}}=\frac{\sum_{t=\lceil k/2\rceil}^{k}\binom{f}{t}\binom{n-f}{k-t}}{\binom{n}{k}}$,
which is the probability that malicious nodes take up the majority of sampled nodes. While this probability is non-trivial even if $f<\frac{n}{2}$, our design can ensure security via the disputing design, with the intuition that as long as the honest nodes take up the majority of global nodes, the re-verification will be more favorable to honest inferencers than dishonest inferencers, so we can design proper incentives such that inferencers will be incentivized to dispute if and only if they are honest but rejected. The detailed analysis is deferred to Appendix E.
@@
Thus, VeriLLM maintains strong probabilistic safety guarantees without global replication. To further reduce risk under targeted committee attacks, the system employs a dynamic verifier-selection mechanism that re-samples verifiers over epochs, ensuring that adversarial control of a fixed subset cannot persist across rounds. Together with a re-verification fallback procedure, this mechanism ensures to detect and reports inconsistencies with overwhelming probability.
5. Protocol Design
@@
Figure 1. System architecture of VeriLLM. The scheduler employs a verifiable random function (VRF) for unbiased node selection and hidden-state sampling. Each node group performs prefill and decoding, commits hidden states, and submits proofs to the blockchain, which verifies results, counts votes, and applies reward or penalty mechanisms to ensure verifiable decentralized inference.
5.1. System Overview
@@
VeriLLM is a publicly verifiable, decentralized inference framework for large language models (LLMs) built atop a blockchain substrate. The system targets three objectives: publicly auditable correctness, low verification overhead compatible with real-time inference, and transparent incentive enforcement via smart contracts. To achieve these goals, VeriLLM consists of three core components: a scheduler, a set of homogeneous node groups, and on-chain verification contracts, which jointly coordinate inference execution, verification, and rewards to ensure correctness and accountability.
@@
  • Scheduler. The scheduler orchestrates task placement, resource allocation, and verification planning. For each inference job, it selects candidate node groups and assigns one Inferencer and multiple Verifiers per group. Although logically centralized, its role assignment and checkpoint sampling are publicly verifiable, as they are derived from a verifiable random function (VRF) (Micali et al., 1999). This design prevents adversaries from predicting verifier roles or biasing selection. The scheduler also commits job metadata and Merkle roots of intermediate states to the blockchain for immutable traceability.
  • Node Groups. Computation is organized into groups of homogeneous GPU workers, each corresponding to identical Transformer layers and execution configurations, consistent with modern tensor- and pipeline-parallel inference (Rajbhandari et al., 2022; Shoeybi et al., 2019). Nodes may act as either Inferencers or Verifiers depending on the scheduler’s assignment. To mitigate role-based adversarial behavior, both roles share an identical execution interface, making inference and verification indistinguishable to the worker during the Prefill phase. This task-type indistinguishability discourages strategic deviation, as misbehavior risks immediate detection.
@@
  • Verification Contracts. Verification is enforced through on-chain smart contracts (Buterin, 2014; Wood, 2014). Nodes stake collateral upon participation, and contracts automatically distribute rewards or penalties based on verification outcomes. To reduce gas costs, VeriLLM adopts a two-tier design: computation and Merkle proof generation occur off-chain, while only compact digests, VRF proofs, and sampled Merkle openings are submitted on-chain. This ensures full public verifiability with on-chain costs sublinear in model size and enables trustless third-party auditing via contract event logs.
Overall, VeriLLM integrates cryptographic randomness, hardware symmetry, and economic incentives within a unified architecture. The scheduler enforces non-manipulable randomness, node groups provide execution indistinguishability, and smart contracts ensure immutable verification and incentive enforcement. Together, these components enable decentralized inference to achieve correctness and efficiency comparable to centralized systems while remaining fully open and publicly verifiable.
End-to-end workflow. (1) A client submits an inference request, after which the scheduler selects node groups and assigns Inferencers and Verifiers using a VRF. (2) The Inferencer executes the forward pass and streams committed hidden states to the scheduler. (3) Verifiers independently reproduce and compare the corresponding hidden states and submit results on-chain. (4) The scheduler samples hidden-state indices to prevent lazy verification, and the smart contract finalizes rewards or penalties based on the on-chain comparison.
5.2. Registration